how many kilograms of tea worth $4.95 a kg should be blended with tea worth $3.87 a kg to make 40kgs of blended tea to sell at $4.35 a kg?
Answer by jorel1380(3719)

x+y=40
4.95x+3.87y=174
3.87x+3.87y=154.8
1.08x=19.2
x=17.78
y=22.22
17.78 kgs of $4.95 tea must be blended with 22.22 kgs of $3.87 tea to make the desired mixture..